Cereal cost 1/3 as much as bacon. Bacon cost 5/4 as much as eggs. Eggs cost what percentage of the cost of cereal?
Answer by Edwin McCravy(20059)

Cereal cost 1/3 as much as bacon.
Then suppose cereal costs $1 and bacon costs $3

Bacon cost 5/4 as much as eggs.
We divide $3 by 5/4.  $3÷(5/4) = $3×(4/5) = $12/5 = $2.40

Eggs cost what percentage of the cost of cereal?
We divide $2.40 by $1.00 and get 2.40 and multiply by 
100 and tack on a % and get 240%.

Edwin